the difference of 2 numbers is 20. the larger is 2 more than 3 times the smaller. find the 2 numbers

Answer by checkley79(3341)

X-Y=20 OR X=20+Y
X=3Y+2
20+Y=3Y+2
Y-3Y=2-20
-2Y=-18
Y=-18/-2
Y=9 ANS. FOR THE SMALLER NUMBER.
X-Y=20
X=20+9
X=29 ANS FOR THE LARGER NUMBER.
PROOF:
29=3*9+2
29=27=2
29=29